OpenStudy (anonymous):

i will give medal and i dont have a calculater so no juging pleaseWhat is the value of the expression? (–1.9 + 16) + (–7.1) A. –25 B. –10.8 C. 7 D. 21.2

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

C. 7

OpenStudy (anonymous):

no its B.

OpenStudy (anonymous):

No, (–1.9 + 16) = 14.1 - 7.1 = 7

OpenStudy (anonymous):

how u did the ? lol

OpenStudy (anonymous):

Chrissy, it's C :)

OpenStudy (anonymous):

ohh it's just that openstudy doesn't recognize some characters or just doesn't understand it.

OpenStudy (anonymous):

like wht?

OpenStudy (anonymous):

idk it was a lessthan and greater than sign last time.

OpenStudy (anonymous):

sorry i was telling some guy to leave me alone because he called me fake

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

It's C.